A store mixes Kenyan coffee worth $10 per kilogram and Venezuelan coffeee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 420-kilogram mixture.
Answer by stanbon(57307)

A store mixes Kenyan coffee worth $10 per kilogram and Venezuelan coffeee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 420-kilogram mixture.
----
Using 2 variables you get:
Equations:
Quantity Eq: K + V = 420 kg
Value Eq:: 10K + 12V = 11*420
-------------------------------------
Multiply thru the Quantity Eq by 10 to get:
10K + 10V = 10*420
-----
Subtract that from the Value Eq. and solve for "V":
2V = 420
V = 210 kg (amt. of Venezuelan cofee needed in the mixture)
----
Since K + V = 420, K = 210 kg (amt. of Kenyan coffee needed in the mix)
================================
Cheers,
Stan H.